Find the unit tangent vector at the point 2,0,pi for a curve which is described by the parametric equations

x=2Sina y=3cosa z=2a


1
Expert's answer
2021-07-18T14:58:31-0400
"\vec v(a)=\vec r'(a)=\langle2\cos a, -3\sin a, 2\rangle"

"||\vec v(a)||=\sqrt{4\cos^2 a+9\sin^2 a+4}"

"=\sqrt{8+5\sin^2 a}"

Point "(2, 0, \pi)"


"a=\dfrac{\pi}{2}"


"\vec v(a)=\vec r'(a)=\langle0, -3, 2\rangle"


"||\vec v(a)||=\sqrt{8+5(1)^2}=\sqrt{13}"

"\vec T(\dfrac{\pi}{2})=\dfrac{-3\vec j+2\vec k}{\sqrt{13}}"

"\vec T(\dfrac{\pi}{2})=\langle0, -\dfrac{3\sqrt{13}}{13}, \dfrac{2\sqrt{13}}{13}\rangle"